OpenStudy (anonymous):

solve the equation by completing the square..2y^2+9=10y

OpenStudy (anonymous):

\[2y ^{2}-10y+9=0\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

still need a hand on this or did you figure it out?

OpenStudy (anonymous):

need a hand please

OpenStudy (anonymous):

so to complete the square, you have to add some constant value to both sides of the equation such that the equation can be written as (ay+b)^2=c so start by moving everything to one side of the equation: \[2y^2-10y+9=0\] Then I'd recommend dividing everything by 2 to simplify things a bit later on. then you have to realize that in order to get the equation into (ay+b)^2 form, the expanded form must be in the form a^2y^2+2ayb+b^2 so: \[y^2-5y+9/2+x=a^2y^2-2aby+b^2\] where x is the constant you will need to add to both sides of the equation. With this you can see that \[a^2=1 \rightarrow a=1\] then solve for b next: \[-5 = -2ab \rightarrow 5/2=b \] \[b^2=25/4=6.25\] finally you need your constant value to match up with this b value. currently you have 2y^2-10y+9=0 and you want b^2 to equal 6.25 or 25/4 to complete the square. To do this add 7/4 or 1.75 from both sides of the equation this leaves you with: \[y^2-5y+25/4=7/4\] Now this is where it all comes together: now the equation is in a perfect square form, so it can be written as \[(y-5/2)^2=7/4\] Finally you get to solve by taking the square root of each side, then adding 5/2 to each side

OpenStudy (anonymous):

So the final answer comes to 7/4?

OpenStudy (anonymous):

not quite. y doesnt equal 7/4; (y-2.5)^2=7/4 to solve for y, which is what you wanted to solve for, take the square root of both sides, so y-2.5=sqrt(7)/sqrt(4) = sqrt(7)/2 then add 2.5 to each side to get y=sqrt(7)/2+2.5 or y=(sqrt(7)+5)/2 unfortunately the problem didn't give the nicest numbers so that's about as clean as the answer's going to get

OpenStudy (anonymous):

That has been my problem, the answer it is giving does not make much sense to me. But well, I guess this is it

OpenStudy (anonymous):

Thanks for your help

OpenStudy (isaiah.feynman):